OpenStudy (anonymous):

a bag contains 7 walnuts, 4 caschews, and 3 pistachio nuts. two of the nuts fall out at random. what is the probability that one is a caschew and the other is a walnut?

OpenStudy (kropot72):

\[P(cashew)=\frac{4}{14}\] \[P(walnut)=\frac{7}{13}\] \[P(cashew\ plus\ walnut)=\frac{4\times7}{14\times13}=you\ can\ calculate\]